(a+4)(a+2) = a^2 +8 ???????????????????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

LHS = (a+4)(a+2) = a(a+2)+4(a+2)
= a^2+2a+4a+8
= a^2+6a+8
≠a^2+8 = RHS
The correct statement is (a+4)(a+2) = a^2+6a+8

Thanks